# Context Matters: Learning Generalizable Rewards via Calibrated Features

### v2

@@ 3.3 Learning calibrated features
We leverage paired comparison queries for learning calibrated features, drawing inspiration from prior work on learning features from human input bobu2022inducing . Unlike traditional comparison approaches that ask users to select the state with the greatest reward, we want to solicit input on how changes in state influence the saliency of an individual feature. We present users with contextual feature queries, where they are shown two states and asked to pick in which state a specific feature is better respected, focusing their attention on that feature alone.
@@
Formally, we collect a set of paired comparisons $\mathcal{D}_{\phi_{i}}=\{(s_{1}^{k},s_{2}^{k},y^{k})\}_{k=1}^{K_{i}}$ for each context-calibrated feature $i\in\{1,\ldots,M\}$, where $K_{i}$ is the number of paired comparisons for feature $i$. Paired comparison labels $y^{k}\in\{0,0.5,1\}$ allow the person to indicate whether the feature is more respected in $s_{1}^{k}\ (y^{k}=1)$, in $s_{2}^{k}\ (y^{k}=0)$, or equivalently across states $(y^{k}=0.5)$. Equivalence queries are crucial for identifying which portions of the state should not affect feature saliency, and contexts in which feature saliency should remain static despite changes in input state value. We split the dataset $\mathcal{D}_{\phi_{i}}$ into two subsets which contain all queries labeled with equivalence, $\mathcal{D}_{\phi_{i}}^{\text{equiv}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\phi_{i}}|y=0.5\}$ and all others, $\mathcal{D}_{\phi_{i}}^{\text{pref}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\phi_{i}}|y\neq 0.5\}$.
@@
Following bobu2022inducing we model human feature comparisons using the Bradley-Terry framework bradley1952rank ,
$P_{\psi_{i}}(s_{1}\succ s_{2})=\frac{\exp(\phi^{\prime}_{\psi_{i}}(s_{1}))}{\exp(\phi^{\prime}_{\psi_{i}}(s_{1}))+\exp(\phi^{\prime}_{\psi_{i}}(s_{2}))}\quad,$ (1)
@@
and utilize a cross-entropy loss over a dataset $D$ of paired comparisons,
$\ell(\psi_{i};D)=\sum_{(s_{1},s_{2},y)\in D}\left[-y\log P_{\psi_{i}}(s_{1}\succ s_{2})-(1-y)\log P_{\psi_{i}}(s_{1}\prec s_{2})\right].$ (2)
@@
$\mathcal{L}(\psi_{i})=\frac{1}{|D_{\phi_{i}}|}\left(\lambda_{\text{FE}}\cdot\ell(\psi_{i};D_{\phi_{i}}^{\text{equiv}})+\ell(\psi_{i};D_{\phi_{i}}^{\text{pref}})\right),$ (3)
where $\lambda_{\text{FE}}$ is a hyperparameter that scales the contribution of state pairs labeled with equivalence. Since calibrated features are learned individually, users can specify whether each feature depends on context or not, eliminating unnecessary data collection for contextually-unaffected features.
3.4 Learning rewards with calibrated features
In this work, we focus on reward learning from preferences, but our framework could also be adapted to demonstrations finn2016gcl , corrections bajcsy2017learning , or any other reward learning approaches. We ask the human reward preference queries, presenting them with two states and asking them to select which state they prefer. We collect a set of paired comparisons $\mathcal{D}_{\theta}=\{(s_{1}^{n},s_{2}^{n},y^{n})\}_{n=1}^{N}$, where $N$ is the number of reward comparisons. We enable equivalence labels such that $y^{n}\in\{0,0.5,1\}$, indicates if $s_{1}^{n}$ is preferred $(y^{n}=1)$, if $s_{2}^{n}$ is preferred $(y^{n}=0)$ or if the states are viewed as equivalent $(y^{n}=0.5)$. Again, paired comparisons are split into subsets which contain all queries labeled with equivalence, $\mathcal{D}_{\theta}^{\text{equiv}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\theta}|y=0.5\}$ and all others, $\mathcal{D}_{\theta}^{\text{pref}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\theta}|y\neq 0.5\}$. We model human reward preferences using the same Bradley-Terry framework (Eq. 1) and employ cross entropy loss (Eq. 2), resulting in the final loss function,
$\mathcal{L}(\psi_{i})=\frac{1}{|D_{\theta}|}\left(\lambda_{\text{RE}}\cdot\ell(\theta;D_{\theta}^{\text{equiv}})+\ell(\psi_{i};D_{\theta}^{\text{pref}})\right),$ (4)
@@ 3.4 Learning rewards with calibrated features
where $\lambda_{\text{RE}}$ is a hyperparameter that scales the contribution of equivalent state pairs.
4 Evaluating impact of calibrated features on reward accuracy
@@
We first evaluate our method using simulated human input to objectively measure reward accuracy against ground truth. Our experiments each consist of two stages: (1) representation learning, and (2) reward learning on these representations. We compare rewards learned on calibrated feature representations versus multi-task baseline representations, investigating whether: (1) explicitly capturing contextual effects using calibrated features yields more accurate rewards than multi-task baselines, and (2) learned calibrated features can be effectively composed to generalize to new test rewards.
@@
Environments. We designed three robotic manipulation environments simulated with PyBullet coumans2016pybullet featuring a Jaco robot in a tabletop setting (Fig. 2). Each environment was constructed to reflect everyday scenarios where human preferences may naturally adapt based on varying environmental conditions and object properties. Each environment contains three task-relevant features (described below and in App. B), which enable rich reward composition further complexified by contextual effects. The continuous state space for each environment is a 23-dimensional vector: 3D end-effector (EE) position and rotation matrix (9 values), 3D positions of three fixed objects (laptop, stove, and human), and two environment-specific contextual elements. For experimental evaluation and simulating human responses, we designate these two state elements as the ground truth contextual factors that influence feature saliency (but don’t assume knowledge of the subset at training). The environments facilitate personal contextual relationships, such as the desire to carry heavier objects closer down to the table, or keep sharper objects farther from the human and pointed away.
@@
Weighted Block. Robot carries a block with variable weight. Features are EE xy-planar stove distance and laptop distance, and EE height above table. Contextual elements are stove heat and block weight.
@@
Cup. Robot carries a cup with variable fullness. Features are EE xy-planar stove distance and laptop distance, and EE cup angle. Contextual elements are stove heat and cup fullness.
@@
Utensil. Robot carries a utensil with variable sharpness. Features are EE xy-planar stove distance and human distance, and utensil point at human. Contextual elements are stove heat and utensil sharpness.
Simulating human response. We simulate human responses by defining ground truth calibrated feature functions and reward models. We define seven ground truth calibrated feature functions (App. C) for the seven unique base features across environments, which serve as examples of desired contextual relationships. Ground truth rewards are linear combinations of calibrated features, representing tradeoffs such as it being more important to keep sharp objects away from a person than to avoid heating up the EE. We simulate human responses to state comparison queries using the Bradley-Terry model of human decision making bradley1952rank . We consider states equivalent if their calculated calibrated feature values or rewards fall within an empirically determined threshold (App. C).
@@
Baselines. Following prior work showing multi-task pretraining to be equally or more effective than meta-RL zhao2022effectiveness , we chose a multi-task baseline with variants differing in the number of training tasks. We evaluate our method against JointPref, the most recent applicable, architecture-agnostic, multi-task IRL method yamada2022task where the representations are learned by training $N$ reward heads on a shared representation. This method pressures the shared representation to compactly capture information relevant across all training tasks. We adapt their approach to use preference queries for reward learning, and test implementations with $N=\{10,25,50\}$ (JointPref-N), in addition to $N=1$, representing a single training preference with uniformly weighted features (SinglePref).
4.1 Experimental design
We compare our method to multi-task baselines in two experiments where context affects one or all reward features. In both, we assess accuracy of 10 held-out test rewards learned from preference queries on calibrated features or multi-task representations. We generate 10k states per environment (80/20 train/test split repeated over 6 seeds). We evaluate reward accuracy as the average win rate over pairwise test state comparisons against ground truth. See all experiment details in App. C.
Experiment 1: Single contextually-affected feature. We first aim to evaluate whether explicitly learning contextual effects as in our proposed approach has an advantage over implicit capture by baseline methods. We do this by investigating settings where a single contextual relationship is being learned, in other words only one of the three reward features is affected by context. Since we operate on human-interpretable features, we assume users can identify this contextually-affected feature and we use all contextual queries to train the associated calibrated feature. Letting $i$ denote the index of the calibrated feature, we assume the remaining calibrated feature functions are identity mappings, i.e., $\phi_{\psi_{j}}^{\prime}(s)=\phi_{j}(s)$ for $j\neq i$, and leave them unmodified. In contrast, the baseline method implicitly learns contextual effects through general reward preference queries. Each representation learning method is trained on 100 pre-training queries (rationale in App. C).
Experiment 2: Multiple contextually-affected features. We next investigate if learned calibrated features can be composed to learn accurate downstream rewards. To do this, we use settings where all three reward features vary with context. While we can compare directly with the baseline in this setting, it is important to note that unlike the baseline our method can also compose with new downstream features – regardless of whether they are context-calibrated or not. In this setting our method no longer has an advantage in targeting all pre-training queries to a single calibrated feature. Each representation learning method is trained on 300 pre-training queries (rationale in App. C), which our method splits evenly between the three features.
Figure 2: Calibrated features learned using simulated human input. Calibrated features reshape base feature given changes in affecting contextual element. Visualization details in App. C
Figure 3: Experiment 1 results (single contextually-affected feature). Each row represents the results for reward capturing contextual effects on each of the three features in an environment.
Figure 4: Experiment 2 results (multiple contextually-affected features). Each figure represents the results for capturing contextual effects on all features in an environment.
4.2 Results
@@
Qualitative analysis of learned calibrated features. Fig. 2 visually compares base features to learned calibrated features, shown across changes in the relevant contextual elements. Results show that learned representations successfully capture meaningful feature shaping functions that appropriately calibrate features based on context. In the first row, the representation adapts the base feature of table distance to capture the effect of object weight, i.e. the heavier the object, the closer to the table it should be carried. The concept of safe carrying is encoded by capturing the relative tradeoff of states under different weights – when no object is carried, all reshaped feature values similar, whereas as the object weight increases there is a sharper contrast in calibrated feature values at higher and lower positions. A similar trend appears across all rows: increasing the stove heat, the sharpness of the utensil, and the cup fullness proportionally affects the stove distance, the human distance, and the laptop distance, respectively.
Reward accuracy with single calibrated feature. Fig. 3 illustrates the accuracy of rewards learned by our method compared to multi-task baselines when a single feature in each environment is contextually-affected. Our method consistently outperforms the best-performing baseline across all experiments, with up to 15% improvement in low-data regimes (5-10 pref queries). See details in App. C. The top two performing baselines in most cases are JointPref (10) and SinglePref models. This is likely because as representations are trained only on 100 queries in this experiment, each reward head for JointPref (25) and JointPref (50) only receive 4 and 2 queries each, respectively. This limited data could prevent them from learning a meaningful representation. Notably, when optimizing only a single calibrated feature, the ground truth reward function appears simple enough that SinglePref model performs relatively well but still falls short of our method’s performance.
@@
Reward accuracy with calibrated feature composition. Results (Fig. 4) demonstrate that our approach continues to outperform the best multi-task baseline across all environments, though with a narrower performance gap in this more complex setting. JointPref (10) and JointPref (25) emerge as the strongest baseline methods across most environments, with SinglePref maintaining competitive performance. This aligns with expectations, as the increased number of pre-training queries provides JointPref methods more data per reward head, enabling better learning of meaningful representations.
Figure 5: Subjective user ratings and model ranking improved with increasing training queries.
5 Evaluating human teaching of calibrated features
Given demonstration of our approach’s benefits in simulated human experiments, we ask: can real humans effectively respond to contextual feature queries to teach their personal contextual preferences? Since user contextual preferences lack ground truth definitions, participants subjectively evaluate whether labeling increasing amounts of queries leads to better contextual preference capture.
5.1 User Study

### v3

@@ 3.3. Learning calibrated features
We leverage paired comparison queries for learning calibrated features, drawing inspiration from prior work on learning features from human input (Bobu et al., 2022). Unlike traditional comparison approaches that ask users to select the highest-reward state, we solicit input on how changes in state influence the saliency of an individual feature. We present users with contextual feature queries, where they are shown two states and asked to pick in which state a specific feature is better respected, focusing their attention on that feature alone.
@@
For each contextually affected feature $\phi_{i}$, we collect a set of $M_{i}$ paired comparisons $\mathcal{D}_{\phi_{i}}=\{(s_{1}^{m},s_{2}^{m},y^{m})\}_{m=1}^{M_{i}}$. Paired comparison labels $y\in\{0,0.5,1\}$ let the user indicate whether the feature is more respected in $s_{1}\ (y=1)$, in $s_{2}\ (y=0)$, or equivalently across states $(y=0.5)$. Equivalence information is crucial for identifying which portions of the state should not affect feature saliency, and contexts in which feature saliency should remain static despite changes in input state value. We split the dataset $\mathcal{D}_{\phi_{i}}$ into two subsets which contain all queries labeled with equivalence, $\mathcal{D}_{\phi_{i}}^{\textnormal{equiv}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\phi_{i}}\mid y=0.5\}$ and all others, $\mathcal{D}_{\phi_{i}}^{\textnormal{pref}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\phi_{i}}\mid y\neq 0.5\}$.
@@
Following (Bobu et al., 2022) we model human feature comparisons using the Bradley-Terry framework (Bradley and Terry, 1952),
(2) $P_{\psi_{i}}(s_{1}\succ s_{2})=\frac{\exp(\phi^{\prime}_{\psi_{i}}(s_{1}))}{\exp(\phi^{\prime}_{\psi_{i}}(s_{1}))+\exp(\phi^{\prime}_{\psi_{i}}(s_{2}))},$
@@
and use a cross-entropy loss over a dataset $\mathcal{D}$ of paired comparisons,
(3) $\ell(\psi_{i};\mathcal{D})=\sum_{(s_{1},s_{2},y)\in\mathcal{D}}\left[-y\log P_{\psi_{i}}(s_{1}\succ s_{2})-(1-y)\log P_{\psi_{i}}(s_{1}\prec s_{2})\right].$
@@
(4) $\mathcal{L}(\psi_{i})=\frac{1}{\lvert\mathcal{D}_{\phi_{i}}\rvert}\left(\lambda_{\text{FE}}\cdot\ell(\psi_{i};\mathcal{D}_{\phi_{i}}^{\text{equiv}})+\ell(\psi_{i};\mathcal{D}_{\phi_{i}}^{\text{pref}})\right),$
@@
where $\lambda_{\text{FE}}$ is a hyperparameter that scales the contribution of equivalence-labeled pairs. Alg. 1 details our learning method.
3.4. Learning rewards with calibrated features
In this work, we learn reward from preferences, but our framework could be adapted to any other reward learning approach, such as demonstrations (Finn et al., 2016) or corrections (Bajcsy et al., 2017). We ask the human reward preference queries, presenting them with two states and asking them to select which state they prefer. We collect a set of paired comparisons $\mathcal{D}_{\theta}=\{(s_{1}^{m},s_{2}^{m},y^{m})\}_{m=1}^{M_{\text{REW}}}$, where $M_{\text{REW}}$ is the number of reward comparisons. We enable equivalence labels such that $y\in\{0,0.5,1\}$, indicates if $s_{1}$ is preferred $(y=1)$, if $s_{2}$ is preferred $(y=0)$ or if the states are viewed as equivalent $(y=0.5)$. Again, paired comparisons are split into subsets which contain all queries labeled with equivalence, $\mathcal{D}_{\theta}^{\text{equiv}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\theta}\mid y=0.5\}$ and all others, $\mathcal{D}_{\theta}^{\text{pref}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\theta}\mid y\neq 0.5\}$. We model human reward preferences with the same Bradley-Terry framework (Eq. 2) and employ cross entropy loss (Eq. 3), resulting in the final loss,
(5) $\mathcal{L}(\theta)=\frac{1}{\lvert\mathcal{D}_{\theta}\rvert}\left(\lambda_{\text{RE}}\cdot\ell(\theta;\mathcal{D}_{\theta}^{\text{equiv}})+\ell(\theta;\mathcal{D}_{\theta}^{\text{pref}})\right),$
where $\lambda_{\text{RE}}$ scales the contribution of equivalence-labeled pairs.
Left panel shows three manipulation environments. The robot holds a separate object in each (block, utensil, or a cup). Right panel shows calibrated feature evaluated at all states in a point cloud above a table top environment. Calibrated features shown for block, stove, utensil, and cup features across a range of the relevant contextual element.
Figure 2. Left: three manipulation environments. Right: calibrated features learned from simulated human input. Learned contextual effects reshape the base feature as the relevant contextual element changes. Point clouds show 8k $E$ locations, normalized over the full range of the relevant contextual element. Irrelevant contextual elements fixed to zero.
. Left panel shows three manipulation environments. The robot holds a separate object in each (block, utensil, or a cup). Right panel shows calibrated feature evaluated at all states in a point cloud above a table top environment. Calibrated features shown for block, stove, utensil, and cup features across a range of the relevant contextual element.
4. Experiments in Simulation
@@
We first evaluate our method using simulated human input to objectively measure reward accuracy against ground truth. Our experiments each consist of two stages: (1) representation learning, and (2) reward learning using learned representations. We compare rewards learned using calibrated features versus multi-task baseline representations, investigating whether: (1) explicitly capturing contextual effects using calibrated features yields more accurate rewards than baselines, and (2) learned calibrated features can be effectively composed to learn new downstream rewards.
4.1. Environments
@@
We designed three robotic manipulation environments simulated with PyBullet (Coumans and Bai, 2016) featuring a Jaco robot in a tabletop setting (Fig. 2). Environments were constructed to reflect everyday scenarios where desired behavior may adapt based on varying environmental conditions and object properties, such as the desire to carry heavier objects closer to the table and farther from electronics, or keep sharper objects farther from a human and pointed away. Preferred adaptations can be personal, reflecting individual differences in traits like caution or comfort. Each environment contains three task-relevant features, which enable rich reward composition further complexified by contextual effects. The continuous state for each environment is 23-dimensional: 3D end-effector (EE) position and rotation matrix (9 values), 3D positions of three fixed objects (laptop, stove, and human), and two environment-specific contextual elements.
Input: Features $\Phi$, $\{M_{i}\}_{i=1}^{K}$ number of queries per feature
for $i\leftarrow 1$ to $\lvert\Phi\rvert$ do
 $c_{i}\leftarrow$ QueryHuman // ask if $\phi_{i}$ contextually affected
 if $c_{i}$ then
    $\mathcal{D}_{\phi_{i}}\leftarrow\{\}$
    for $j\leftarrow 1$ to $M_{i}$ do
       $(s_{1},s_{2})\leftarrow\textsc{Uniform}(\mathcal{S})$ // sample (any strategy)
       $y\leftarrow$ ContextualQuery($\phi_{i},s_{1},s_{2}$) // ask human
       $\mathcal{D}_{\phi_{i}}\leftarrow\mathcal{D}_{\phi_{i}}\cup(s_{1},s_{2},y)$
      end for
    $\phi^{c}_{\psi_{i}}(s)\leftarrow\text{train\_calib\_feat}(\mathcal{D}_{\phi_{i}})$ // via Eq. 4
    $\phi^{c}_{i}(s)\leftarrow\phi^{c}_{\psi_{i}}(s)$
   end if
 else
    $\phi_{i}^{c}(s)\leftarrow\phi_{i}(s)$
   end if
end for
Algorithm 1 Learning Calibrated Features
For evaluation and simulating human responses, we designate these two state elements as the ground truth contextual factors that influence feature saliency (but don’t assume knowledge of the subset at training). The three environments are:
@@
Weighted Block. The robot carries a block with variable weight. Features are: (1) stove, EE xy-planar stove distance, (2) laptop, EE xy-planar laptop distance, and (3) table, EE height above table. Contextual elements are stove heat and block weight.
@@
Cup. The robot carries a cup with variable fullness. Features are: (1) stove, (2) laptop, and (3) cup angle, defined as the projection of the EE’s x-axis (cup’s upright axis) onto the world vertical axis. Contextual elements are stove heat and cup fullness.
@@
Utensil. The robot carries a utensil with variable sharpness. Features are: (1) stove, (2) human, EE xy-planar human distance, and (3) point, defined as the projection of the end-effector’s x-axis (utensil’s pointing direction) onto the unit vector pointing from the EE to the human. Contextual elements are stove heat and utensil sharpness.
For more details on base feature implementation, see App. B.1.
[…]
4.2. Experimental Design
Line graph showing reward accuracy against number of reward preference queries. Our method shows higher performance at all non-zero query counts, outperforming all baseline methods. The gap between our method and the best baseline is most pronounced in low data regimes (5-10 preference queries) and closes toward the higher query counts (50 queries).
Figure 3. Exp. 1 & 2 reward accuracy results. Results shown for each environment (rows) when learning rewards that capture contextual effects on one of the three environment features (columns 1-3) or all three environment features (column 4).
Line graph showing reward accuracy against number of reward preference queries. Our method shows higher performance at all non-zero query counts, outperforming all baseline methods. The gap between our method and the best baseline is most pronounced in low data regimes (5-10 preference queries) and closes toward the higher query counts (50 queries).
Bar graph showing induced trajectory reward results for each environment for each single calibrated feature and all calibrated features. Our method outperforms all baseline methods across all experiments.
Figure 4. Exp 1 & 2 induced trajectory reward results, showing ground truth reward of trajectories selected by learned rewards. Results shown for each environment for rewards in which a single or all three environment features are contextually-affected.
Bar graph showing induced trajectory reward results for each environment for each single calibrated feature and all calibrated features. Our method outperforms all baseline methods across all experiments.
Manipulated Variables. We compare four representation learning methods when either a single feature or all features are contextually affected. Each method is allotted a budget of $M_{\text{REP}}$ queries. See App. A for implementation and training details.
(1) Calibrated Features. We assume access to the set of contextually-affected features and train a calibrated feature for each, evenly dividing $M_{\text{REP}}$ contextual feature queries across them.
@@ 4.2. Experimental Design
(2) JointPref. Following prior work showing multi-task pre-training for representation learning to be equally or more effective than meta-RL (Zhao et al., 2022), we compare against JointPref, a multi-task baseline (Yamada et al., 2022) where representations are learned by pre-training $N$ reward heads with shared initial layers. We adapt their approach to use $M_{\text{REP}}/N$ preference queries to train each reward head, testing implementations with $N=\{10,25,50\}$ (JointPref-N). Model input is the vector of base features concatenated with the state. Shared layers compress the input to a 7-dimension embedding.
(3) SinglePref. We implement JointPref with $N=1$, where the sole training preference uniformly weights features.
(4) RandomInit. We include a randomly initialized representation with the same architecture as the multi-task baseline shared layers. This offers a lower bound for the benefits of pre-training.
Box and whisker plots showing user Likert Scale responses and Ranking responses. A separate box is shown for each feature for each number of training queries (0, 50, 25, and 100). Likert scale scores increase as the number of training queries increases, while rankings decrease with increased training queries (given lower is better).
[…]
Box and whisker plots showing user Likert Scale responses and Ranking responses. A separate box is shown for each feature for each number of training queries (0, 50, 25, and 100). Likert scale scores increase as the number of training queries increases, while rankings decrease with increased training queries (given lower is better).
Metrics. We evaluate the quality of the learned representations by learning 10 held-out test rewards on top of each representation type. We test both freezing and fine-tuning representations during reward learning, reporting the stronger variant for each method (our method with freezing, baselines with fine-tuning; see App. C.3). We compare the accuracy of rewards learned with varying amounts of randomly sampled preference queries per test reward, $M_{\text{REW}}$. To ensure evaluation on separate tests states from those the robot is trained on, we generate 10k possible states per environment which we split into an 80/20 train/test distribution (repeated over 6 seeds). To generate training queries, we randomly sample $M_{\text{REP}}$ representation learning queries and $M_{\text{REW}}$ reward learning queries from the distribution of 8k possible train states. Reward accuracy is then measured as the average win rate across all test rewards evaluated over 1k state pairs randomly sampled from the test states. We additionally evaluate the quality of trajectories induced by rewards learned with a fixed low-data preference query budget ($M_{\text{REW}}=10$). Across 100 generated trajectory sets (10 contexts × 10 start–goal pairs), each containing 100 trajectories, we compute the ground-truth reward (normalized within each set) of the trajectory ranked the highest by the learned reward. We report this value averaged over all trajectory sets and all test rewards.
Exp. 1: Single contextually affected feature. We first test whether explicitly learning contextual effects with calibrated features provides an advantage over implicit capture by baselines. We examine cases where only one of the three reward features is affected by context. Each method is trained on $M_{\text{REP}}=100$ pre-training queries, chosen based on model convergence (see App. C.2 for an in-depth analysis of convergence properties of our method and baselines).
Exp. 2: Multiple contextually affected features. We next investigate if learned calibrated features can be composed to learn accurate downstream rewards using scenarios where all three reward features vary with context. We accordingly scale the number of pre-training queries to $M_{\text{REP}}=300$.
4.3. Results
@@
Qualitative analysis of learned calibrated features. Fig. 2 compares base features to learned calibrated features across changes in the relevant contextual element. Results show that learned representations successfully capture meaningful feature shaping functions that appropriately calibrate features based on context. In the first row, the representation adapts the base table feature to capture the effect of object weight, i.e., the heavier the object, the closer to the table it should be carried. The reshaped feature values encode the concept of safe carrying–values are almost uniform when no object is carried, but increasing object weight produces a sharper contrast in values at higher and lower positions. A similar trend appears across all rows: increasing the stove heat, utensil sharpness, and cup fullness affect the stove, human, and laptop features, respectively.
Exp. 1: Single calibrated feature. Fig. 3 columns 1-3 show learned reward accuracy when a single feature is contextually affected. Our method consistently outperforms the best-performing baseline across all experiments, with up to 15% improvement in low-data regimes (5-10 pref queries). Fig. 4 similarly shows trajectories induced by rewards learned using our method outperform the best-performing baseline when a single feature is calibrated. The best-performing baselines across metrics are often JointPref (10) and SinglePref models, although their performance is generally similar to random representation initialization. This indicates that in low-data regimes, implicit representation learning struggles to capture useful structure. This is unsurprising, given SinglePref lacks the signal required to distinguish task-specific reward from general structure. Meanwhile, multi-task variants spread 100 representation queries across all rewards (10 queries per reward for JointPref (10), 4 for JointPref (25) and 2 for JointPref (50)), leaving too little data to distinguish shared contextual effects from task-specific rewards.
@@
Exp. 2: Calibrated feature composition. Results (Fig. 3 &  4) demonstrate that our approach continues to outperform the best baseline across metrics, though with a narrower performance gap in this more complex setting. With 300 pre-training queries, we now see all baselines generally outperform random initialization. This aligns with expectations since a larger pre-training budget gives multi-task methods more data per reward head, enabling better use of pre-training reward signals to learn meaningful representations.
5. User Study
Given the demonstrated benefits of our approach in simulation, we ask: can real humans label contextual feature queries to teach their personal contextual preferences? To investigate, we had participants label contextual feature queries and compare calibrated feature representations trained with varying query amounts, revealing whether increasing labeled input improves preference capture.